The Dallas-Fort Worth International airport ranked 3 in major U.S. airports in on-time arrivals in the last quarter of 2008 with an 82.9% on-time rate. Choose 5 arrivals at random and find the probability that at least 1 was not on time. Please round the final answer to 2 or 3 decimal places. Source: www.census.gov
The Sky Harbor International airport ranked 9 in major U.S. airports in on-time arrivals in the last quarter of 2008 with an 81.6 % on-time rate. Choose 5 arrivals at random and find the probability that at least 1 was not on time. Please round the final answer to 2 or 3 decimal places. Source: www.census.gov P(at least 1 was not on time)
Answer #1

1. P(a flight is on time) = 0.892

Number of flight arrivals = 5

P(at least 1 was not on time) = 1 - P(all were on time)

= 1 - 0.8295

= 0.608

2. P(a flight is on time) = 0.816

Number of arrivals = 5

P(at least 1 was not on time) = 1 - P(all were on time)

= 0.638